# Conditional gene genealogies given the population pedigree for a diploid Moran model with selfing

### v1

@@ 7 Discussion
In this work, we have assumed that the pedigree of a population of constant size $N$ is the outcome of a random process of reproduction in which offspring are produced by self-fertilization with probability $\alpha_{N}$. By conditioning on the pedigree and considering pairwise times to common ancestry, we found three different coalescent models in the limit $N\to\infty$, when time is measured in units of $N$ generations (which is $N^{2}$ time-steps in the Moran model). The deciding factor is $N(1-\alpha_{N})$, which may be interpreted either as the rate of outcrossing events along an ancestral lineage or as the expected number of outcrossed offspring in the population each generation. The three models are ‘negligible outcrossing’ which applies when $N(1-\alpha_{N})\to 0$, ‘limited outcrossing’ which applies when $N(1-\alpha_{N})\to\lambda\in(0,\infty)$, and ‘partial selfing’ which applies when $N(1-\alpha_{N})\to\infty$. Negligible outcrossing and limited outcrossing both require $\alpha_{N}\to 1$, while partial selfing includes this as a special case.
@@
Previous population-genetic analyses of selfing populations have assumed a fixed selfing rate, $\alpha_{N}=\alpha$ in our notation, and have implicitly averaged over pedigrees. When this averaging is done, just one coalescent model for all $\alpha\in[0,1]$ emerges in the limit, specifically the coalescent model with partial selfing described by Nordborg and Donnelly (1997) and Möhle (1998). When coalescence is conditioned on the pedigree and $\alpha$ is fixed, our very similar partial-selfing model is obtained but only for $\alpha\in[0,1)$. Assuming a fixed selfing rate of $\alpha=1$ leads to an entirely different model, namely negligible outcrossing.
@@
A more detailed look at how $\alpha_{N}$ approaches $1$ reveals the critical case of limited outcrossing in between negligible outcrossing and partial selfing and characterized by the parameter $N(1-\alpha_{N})\to\lambda\in(0,\infty)$. In this regime, the parts of the population pedigree directly ancestral to the sampled individuals can be replaced by a random graph like the ones previously used to model recombination (Griffiths, 1991; Griffiths and Marjoram, 1997) and selection (Krone and Neuhauser, 1997; Neuhauser and Krone, 1997). It is distinguished from these only by what the splitting events in the graph represent and how these are treated in modeling gene genealogies. Each splitting event in the graph corresponds to an outcrossing event in the genealogical ancestry of the sampled individuals. Ancestral lineages tracing back to such a node in the graph make a $50$:$50$ choice of which branch they follow. The rate at which ancestral lineages split is precisely our parameter $\lambda$.
@@
As with the well known Wright-Fisher diffusion and the corresponding standard neutral coalescent model, these new models of coalescence conditional on the pedigree are meant as robust approximations for large populations. We expect, for example, that the Wright-Fisher model with selfing used to produce Figure 1 will have the same three limiting cases we found for the Moran model with selfing. Which of the three is most applicable for a given (necessarily finite) population will depend on $N$ and $\alpha_{N}$. The special case of negligible outcrossing involves the strong predictions of identical gene genealogies across all loci in the genome and zero heterozygosity, and so may be of limited utility. The critical case of limited outcrossing allows gene genealogies to differ across the genome but also predicts zero heterozygosity. Still it is clear from the jumps in probability in Figure 1 that if $N(1-\alpha_{N})$ is not large, insights from the limited-outcrossing model are helpful even when the value of $\alpha_{N}$ would not prevent heterozygosity. On the other hand, both Figure 1 and Figure 8 suggest that if $N(1-\alpha_{N})$ is greater than about $100$, partial-selfing is a better description than limited-outcrossing, even if the outcrossing rate $1-\alpha_{N}$ is quite small.
@@
We chose to use the name “partial selfing" here because in this case the coalescent process conditional on the pedigree is very similar to the previously described partial-selfing model which did not condition on the pedigree (Nordborg and Donnelly, 1997; Möhle, 1998). Our model may be considered an update of the previous one. In both, times to common ancestry for “diff” samples follow the Kingman coalescent process with effective population size $N_{e}=(2-s)N/2$. In our formulation, partial selfing is obtained when $N(1-\alpha_{N})\to\infty$. Similarly we may recall from Figure 8 that as $\lambda\to\infty$ the predictions for “diff” samples under limited outcrossing converge on those for partial selfing. The reason that pedigrees do not constrain “diff” coalescence times under the partial-selfing model is that outcrossing events dominate the ancestry of the population. The same underlying phenomena are at work here as in the standard models of randomly mating populations, where the fact that individuals have two parents with high probability strongly influences the pedigree (Chang, 1999; Derrida et al., 1999; Coron and Le Jan, 2022) resulting in a fast mixing time of ancestral genetic processes given the pedigree (Barton and Etheridge, 2011). Because of this, the conditional and unconditional coalescent processes for these standard population models converge to the same Kingman coalescent process (Tyukin, 2015; Diamantidis et al., 2024).
@@
For the purposes of discussion, in what follows we take for granted that our conjecture about the coalescent process of a “diff” sample of size $n>2$ in Section 4.2 holds for partial selfing, namely that it is given by a Kingman coalescent process. We further conjecture that a sample of size $n$ containing $2m$ gene copies together as pairs in $m$ individuals and $n-2m$ gene copies each in different individuals, cf. Nordborg and Donnelly (1997), will experience a random number of effectively instantaneous coalescent events, $X_{i}$, now explicitly at locus $i$, then the remaining $n-X_{i}$ lineages will enter the Kingman coalescent process. Each of the $m$ individuals will have its own number of selfing generations before its first outcrossing event, given by independent draws of the random variable $U$. If these outcomes are $\{k_{1},k_{2},\ldots,k_{m}\}$, then $X_{i}\sim\sum_{r=1}^{m}\textrm{Bernoulli}\left({1-2^{-k_{r}}}\right)$, as in (3) and (4). An unlinked locus, $j\neq i$, has the same $\{k_{1},k_{2},\ldots,k_{m}\}$ but a conditionally independent value of $X_{j}$ and an independent realization of the Kingman coalescent process.
@@
The coalescent with partial selfing conditional on the pedigree accurately predicts the variation in the level of inbreeding among individuals which Haldane (1949), Bennett and Binet (1956) and Kimura (1958) highlighted at the population level, and Weir and Cockerham (1973) developed into a theory of identity disequilibrium; recall (1) and (4). But like all coalescent models, the focus is now on the sample. If we write $Y_{i,r}$ for the indicator random variable of coalescence in the recent ancestry of sampled individual $r\in\{1,2,\ldots,m\}$ for the sample of the previous paragraph, then $E[Y_{i,r}]=s/(2-s)$ as in (3) and $\textrm{Var}[Y_{i,r}]$ is given by (1). If we were able to observe the outcomes $y_{i,r}$ for $r\in\{1,2,\ldots,m\}$, then the mean among loci $\bar{y}_{r}$ would be an estimate of $1-2^{-k_{r}}$ and the sample variance of $\bar{y}_{r}$ among individuals would be an estimate of the identity disequilibrium (1).
Note that our partial-selfing model permits another axis of sampling to be considered, namely the sampling of pedigrees. For example, one may imagine a number of pedigrees independently generated under the model and a single individual sampled from each. The inbreeding coefficient $F=s/(2-s)$ and the identity disequilibrium (1) predict the mean and variance of inbreeding among these individuals.
@@
Identity by descent (i.e. the outcomes $y_{i,r}$ above) cannot be observed directly but only inferred based on patterns of identity in state (Pollak, 1987; Rousset, 2002; Uyenoyama, 2024). A number of methods have been developed and applied to estimate inbreeding coefficients. Some have made use of $U$ in (2) or similar distributions to estimate average inbreeding coefficients of populations (Enjalbert and David, 2000; David et al., 2007; McClure and Whitlock, 2012). Others have employed identity disequilibria (Ritland and Jain, 1981; Ritland, 1996; Sweigart et al., 1999; Yang, 2002; Jarne and David, 2008; Szulkin et al., 2010; Vieira et al., 2013) or runs of homozygosity (Franklin, 1977; Yi et al., 2022; Zeitler and Gilbert, 2024) to estimate individual inbreeding coefficients. Still others have taken up the specific problem of estimating the number of generations back to the most recent outcrossing event for each individual together with the selfing rate (Gao et al., 2007; Wang et al., 2012; Redelings et al., 2015).
@@
The existence of the limited-outcrossing coalescent model, with its simple structure of the random graph, can aid in the interpretation of multi-locus genetic data even if its extreme prediction of zero heterozygosity of individuals is not met. Each branching event in the ancestral graph corresponds to an outcrossing event between ancestral individuals in the pedigree of the sample, in the case that these are rare, i.e. with rate proportional to $1/N$ in the model. Forward in time and to the extent that the two ancestral individuals differ genetically, each such event would create a potentially novel recombinant inbred line. The partial-selfing model approximates this when the selfing rate is high, except that some of these new lines may not become fully inbred.
@@
It would not be as straightforward to incorporate selection or population structure because these can have their own effects on the pedigree. Some simulations demonstrating this can be found in Kuo and Avise (2008), Wakeley et al. (2016) and Wilton et al. (2017). It is an open question how recombination and coalescence operate conditional on the population pedigree. We have dealt with free recombination here. Theorem 4.3 and Theorem 4.4 are about the distribution of conditionally independent coalescence times given the pedigree. For the unconditional or pedigree-averaged case, Kogan et al. (2023) recently considered the full range of recombination rates and how these might be assumed to scale with $N$ in an analysis of correlations of coalescence times in samples of size two at two loci. It may be that different results for recombination and coalescence will be obtained in the conditional case depending on how recombination scales with $N$.

### v2

@@ 6 Discussion
In this work, we have assumed that the pedigree of a population of constant size $N$ is the outcome of a random process of reproduction in which offspring are produced by self-fertilization with probability $\alpha_{N}$. By conditioning on the pedigree and considering pairwise times to common ancestry, we found three different coalescent models in the limit $N\to\infty$, when time is measured in units of $N$ generations (which is $N^{2}$ time-steps) in the Moran model. The deciding factor is $N(1-\alpha_{N})$, which may be interpreted either as the rate of outcrossing events along an ancestral lineage or as the expected number of outcrossed offspring in the population each generation. The three models are ‘negligible outcrossing’ which applies when $N(1-\alpha_{N})\to 0$, ‘limited outcrossing’ which applies when $N(1-\alpha_{N})\to\lambda\in(0,\infty)$, and ‘partial selfing’ which applies when $N(1-\alpha_{N})\to\infty$. Negligible outcrossing and limited outcrossing both require $\alpha_{N}\to 1$, while partial selfing includes this as a special case.
@@
Previous population-genetic analyses of selfing populations have assumed a fixed selfing probability, $\alpha_{N}=\alpha$ in our notation, and have implicitly averaged over pedigrees. When this averaging is done, just one coalescent model for all $\alpha\in[0,1]$ emerges in the limit, specifically the coalescent model with partial selfing described by Nordborg and Donnelly (1997) and Möhle (1998). When coalescence is conditioned on the pedigree and $\alpha$ is fixed, our very similar partial-selfing model is obtained but only for $\alpha\in[0,1)$ when $\alpha$ is fixed. Assuming a fixed selfing probability of $\alpha=1$ leads to an entirely different model, namely negligible outcrossing.
@@
A more detailed look at how $\alpha_{N}$ approaches $1$ reveals the critical case of limited outcrossing in between negligible outcrossing and partial selfing, and characterized by the parameter $N(1-\alpha_{N})\to\lambda\in(0,\infty)$. In this regime, the parts of the population pedigree directly ancestral to the sampled individuals can be replaced by a random graph like the ones previously used to model recombination (Griffiths, 1991; Griffiths and Marjoram, 1997) and selection (Krone and Neuhauser, 1997; Neuhauser and Krone, 1997). It is distinguished from these only by what the splitting events in the graph represent and how these are treated in modeling gene genealogies. Each splitting event in the graph corresponds to an outcrossing event in the genealogical ancestry of the sampled individuals. Ancestral lineages tracing back to such a node in the graph make a $50$:$50$ choice of which branch they follow. The rate at which ancestral lineages split is precisely our parameter $\lambda$.
@@
As with the well known Wright-Fisher diffusion and the corresponding standard neutral coalescent model, these new models of coalescence conditional on the pedigree are meant as robust approximations for large populations. We expect, for example, that the Wright-Fisher model with selfing used to produce Figure 1 will have the same three limiting cases we found for the Moran model with selfing. Which of the three is most applicable for a given (necessarily finite) population will depend on $N$ and $\alpha_{N}$. The special case of negligible outcrossing involves the strong predictions of identical gene genealogies across all loci in the genome and zero heterozygosity, and so may be of limited utility. The critical case of limited outcrossing allows gene genealogies to differ across the genome but also predicts zero heterozygosity. Still it is clear from the jumps in probability in Figure 1 that if $N(1-\alpha_{N})$ is not large, insights from the limited-outcrossing model are helpful even when the value of $\alpha_{N}$ would not prevent heterozygosity. On the other hand, both Figure 1 and Figure 7 suggest that if $N(1-\alpha_{N})$ is greater than about $100$, partial-selfing is a better description than limited-outcrossing, even if the outcrossing probability $1-\alpha_{N}$ is quite small.
@@
We chose to use the name “partial selfing” here because in this case the coalescent process conditional on the pedigree is very similar to the previously described partial-selfing model which did not condition on the pedigree (Nordborg and Donnelly, 1997; Möhle, 1998). Our model may be considered an update of the previous one. In both, times to common ancestry for “diff” samples follow the Kingman coalescent process with effective population size $N_{e}=(2-s)N/2$. In our formulation, partial selfing is obtained when $N(1-\alpha_{N})\to\infty$. Similarly we may recall from Figure 7 that as $\lambda\to\infty$ the predictions for “diff” samples under limited outcrossing converge on those for partial selfing. The reason that pedigrees do not constrain “diff” coalescence times under the partial-selfing model is that outcrossing events dominate the ancestry of the population. The same underlying phenomena are at work here as in the standard models of randomly mating populations, where the fact that individuals have two parents with high probability strongly influences the pedigree (Chang, 1999; Derrida et al., 1999; Coron and Le Jan, 2022) resulting in a fast mixing time of ancestral genetic processes given the pedigree (Barton and Etheridge, 2011). Because of this, the conditional and unconditional coalescent processes for these standard population models converge to the same Kingman coalescent process (Tyukin, 2015; Diamantidis et al., 2024).
@@
Let us take for granted the conjecture we made in Section 4.1 concerning a sample of size $n$ with $2m$ gene copies together as pairs in $m$ individuals. At a locus $i$ in such a sample, there will be a random number of effectively instantaneous coalescent events $X_{i}$, then the remaining $n-X_{i}$ lineages will enter the Kingman coalescent process. If $\{k_{1},k_{2},\ldots,k_{m}\}$ are the realized numbers of selfing generations in the ancestries of the $m$ individuals, then $X_{i}\sim\sum_{r=1}^{m}\textrm{Bernoulli}\left({1-2^{-k_{r}}}\right)$. An unlinked locus, $j\neq i$, has the same $\{k_{1},k_{2},\ldots,k_{m}\}$ but a conditionally independent value of $X_{j}$ and an independent realization of the Kingman coalescent process.
@@
This coalescent model with partial selfing conditional on the pedigree accurately predicts the variation in the level of inbreeding among individuals which Weir and Cockerham (1973) developed into a theory of identity disequilibrium; recall (1) and (4). Like all coalescent models, the focus is now on the sample rather than the population. If we write $Y_{i,r}$ for the indicator random variable of coalescence in the recent ancestry of sampled individual $r\in\{1,2,\ldots,m\}$, then $E[Y_{i,r}]=s/(2-s)$ as in (3) and $\textrm{Var}[Y_{i,r}]$ is given by (1) or (4). If we were able to observe the outcomes $y_{i,r}$ for $r\in\{1,2,\ldots,m\}$, then the mean among loci $\bar{y}_{r}$ would be an estimate of $1-2^{-k_{r}}$ and the sample variance of $\bar{y}_{r}$ among individuals would be an estimate of the identity disequilibrium (1).
@@
Identity by descent cannot be observed directly but only inferred based on patterns of identity in state (Pollak, 1987; Rousset, 2002; Uyenoyama, 2024). A number of methods have been developed and applied to estimate inbreeding coefficients or selfing probabilities. Some have made use of $U$ in (2) or similar distributions to estimate selfing probabilities of populations (Enjalbert and David, 2000; David et al., 2007; McClure and Whitlock, 2012). Others have employed identity disequilibria (Ritland and Jain, 1981; Ritland, 1996; Sweigart et al., 1999; Yang, 2002; Jarne and David, 2008; Szulkin et al., 2010; Vieira et al., 2013) or runs of homozygosity (Franklin, 1977; Yi et al., 2022; Zeitler and Gilbert, 2024) to estimate individual inbreeding coefficients due to selfing. Others have taken up the problem of estimating the number of generations back to the most recent outcrossing event for each individual together with the selfing probability (Gao et al., 2007; Wang et al., 2012; Redelings et al., 2015). Our results for partial selfing establish a coalescent framework for these methods, and provide for new coalescent-based estimators like the one described in Nordborg and Donnelly (1997, Eq. 20) but incorporating (2) and $\{k_{1},k_{2},\ldots,k_{m}\}$ above in place of just $F=s/(2-s)$.
@@
The existence of the limited-outcrossing coalescent model, with its simple structure of the random graph, can aid in the interpretation of multi-locus genetic data even if its extreme prediction of zero heterozygosity of individuals is not met. Each branching event in the ancestral graph corresponds to an outcrossing event between ancestral individuals in the pedigree of the sample, in the case that these are rare, i.e. with probability proportional to $1/N$ in the model. Forward in time and to the extent that the two ancestral individuals differ genetically, each such event would create a potentially novel recombinant inbred line. The partial-selfing model approximates this when the selfing probability is high, except that some of these new lines may not become fully inbred.
@@
It would not be as straightforward to incorporate selection or population structure because these can have their own effects on the pedigree. Some simulations demonstrating this can be found in Kuo and Avise (2008), Wakeley et al. (2016) and Wilton et al. (2017). It is an open question how recombination and coalescence operate conditional on the population pedigree. We have dealt mostly with free recombination here. In particular, Theorem 4.1 and Theorem 4.5 are about the distribution of conditionally independent coalescence times given the pedigree. In the case of limited outcrossing, we were able to include an analysis of two loci with recombination which gives covariances of coalescence times consistent with the discrete, exact results for the unconditional or pedigree-averaged case in Kogan et al. (2023).